HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

IFLEX-AGILE: A COMMERCIAL CW-OPO FOR SPECTROSCOPIC APPLICATIONS

B. PILLES, D. WITSCH and \underline{F. MUELLER} , EXCELITAS TECHNOLOGIES, 85622 Feldkirchen (Munich), Germany, www.excelitas.com

Excelitas Technologies Corp. is a photonics technology leader focused on delivering innovative, high-performance, market-driven solutions to meet the lighting, optronics, detection and optical technology needs of our OEM customers.

With the Excelitas brand LINOS being the first company worldwide ever offering cw-OPOs commercially since 1999, the history is continued with the present iFLEX-Agile series. Ideal for a variety of applications in mass spectrometry, material testing, trace-gas sensing, spectroscopy, and metrology, this compact instrument allows finetuning of specific wavelengths from NIR to MIR, having high-quality output beams of high intensity. Fully controlled via USB computer interface, iFLEX-Agile enables rapid and reproducible settings of any desired wavelength from 1480 up to 1950nm and from 2400 up to 3800nm with output powers exceeding 1W in its basic configuration, and without any need to change optics or modules. iFLEX-Agile emission linewidths can vary from 200GHz to below 1MHz depending on its configuration. The modular concept allows modification of wavelength tuning range, emission linewidths and output power levels to cost-efficiently meet the custom requirements for specific applications.

The authors acknowledge the support by Claudia Gabler, Georg Koch (technical design) and Tatjana Brill (marketing).

Link to product: https://www.excelitas.com/product/iflex-agile-cw-optical-parametric-oscillator

p-number: p225  Submitted on Sun, 18 Jun 2023 19:53:32 +0200